OpenStudy (anonymous):

can someone help me with this problem please :o What are the zeros of the polynomial function: f(x) = (x - 4)(2x + 1)(x + 7)?

OpenStudy (nenadmatematika):

4,-1/2 and -7

OpenStudy (anonymous):

thanks

OpenStudy (nenadmatematika):

you're welcome :D

OpenStudy (anonymous):

simple: (x - 4)(2x + 1)(x + 7) x-4=0 # 2x+1=0 # x+7=0; x=4 # x=-1/2 # x=-7;